OpenStudy (anonymous):

a. Suppose a market research company finds that at a price of p = $20, they would sell x = 42 tiles each month. If they lower the price to p = $10, then more people would purchase the tile, and they can expect to sell x = 52 tiles in a month’s time. Find the equation of the line for the demand equation. Write your answer in the form p = mx + b. Hint: Write an equation using two points in the form (x,p).

OpenStudy (dumbcow):

2 points are (42,20) (52,10) m = (10-20)/(52-42) = -10/10 = -1 20 = (-1)(42) +b b=62 p =-x +62

OpenStudy (anonymous):

b. Substitute the result you found from part a. into the equation R = xp to find the revenue equation. Provide your answer in simplified form.

OpenStudy (dumbcow):

R=x(-x+62) = -x^2 +62x